### v1

# SpiralMap: A Python library of the Milky Way’s spiral arms

SpiralMap: A Python library of the Milky Way’s spiral arms
Abhay Kumar Prusty Indian Institute of Science Education and Research Kolkata, Mohanpur 741246, West Bengal, India [ Shourya Khanna INAF - Osservatorio Astrofisico di Torino, via Osservatorio 20, 10025 Pino Torinese (TO), Italy [
Galaxy structure, Spiral arms, Milky Way, Galactic dynamics
show]akp22ms087@iiserkol.ac.in
]shourya.khanna@inaf.it
Statement of need
Mapping the structure of the Galaxy has been an ongoing endeavour for several decades, thanks to which we have come to piece together the components it is made up of such as the discs/bulge/halo and so on (Bland-Hawthorn & Gerhard, 2016). Of particular interest is also the signature of non-axisymmetry in the disc, principally present in the form of the central Galactic bar, the warp, and various spiral like features fanning across a large portion of the disc. Over the years, various groups have surveyed the Galaxy across wavelengths (Radio to optical) and have deduced the rich variety of spiral structure present in the disc. While some of the papers in literature provide machine readable data to trace out the spiral arms in their model, often this can be a cumbersome exercise for a user simply interested in extracting the coordinates and/or overplotting the spiral arms on another plot of interest, such as while comparing the locations of the arms to features in the velocity field (Khanna et al., 2023; Poggio et al., 2024, etc.). We show an example science application in Figure 4.
With SpiralMap we present a library of the major spiral arm models (and maps) of the Galaxy. The package is written in Python, and allows the user to both extract the 2D trace and overplot the spiral arms in cartesian/polar coordinates in both Heliocentric (HC) and Galactocentric (GC) frames. A summary of the models currently included is provided in Table 1, where we have tried to include [^1] Fair practice: If you make use of our package, please ALSO cite the individual model papers. A bib file is provided in readthedocs. models from across the electromagnetic spectrum, and based on various tracers (gas/stars etc.). Other models can easily be included upon request. In the near future, we anticipate the availability of 3D spiral arm traces for the Galaxy in literature which can also be included in SpiralMap. A few example plots that can be generated using the package are included below. For full details, we point to the following links: a) Documentation: readthedocs, b) Demonstration: Jupyter notebook, c) GitHub repository.
[…]
Table 1: Basic summary of the spiral arm models included in SpiralMap.
Model Description
Taylor_Cordes_1992 Model based on HII (Taylor & Cordes, 1993).
Drimmel_NIR_2000 Model based on Galactic plane emission in the NIR (Drimmel, 2000).
Levine_2006 Model based on HI (21 cm) (Levine et al., 2006).
Hou_Han_2014 Logarithmic spiral model based on HII/ GMC/methanol Maser observations (Hou & Han, 2014).
Reid_2019 Model based on high-precision parallax measurements of MASERS (Reid et al., 2019).
Poggio_2021 Map based on Upper Main sequence stars (Poggio et al., 2021)
Gaia_2022 Map based on OB stars (Gaia Collaboration et al., 2023)
Drimmel_Ceph_2024 Model based on Cepheid variables (Drimmel et al., 2024).
Figure 1: Cartesian projection of the Drimmel_Ceph_2024 model shown for a particular arm (Sag-Car). We show this arm in HC (a), HC with a polar grid in the background (b), and in GC frame with a polar grid in the background (c).
Figure 2: Cartesian projections of multiple models plotted together with a polar grid in the background. We show the Taylor_Cordes_1992 & Poggio_2021 models in HC (a) and GC (b) frames, and similarly, the Drimmel_NIR_2000 & Poggio_2021 models in HC (c) and GC (d) frames.
Figure 3: Polar projections of multiple models plotted together. We show the Taylor_Cordes_1992 & Poggio_2021 models in HC (a) and GC (b) frames, and similarly the Drimmel_NIR_2000 & Poggio_2021 models in HC (c) and GC (d) frames.
1 Available models
In the current version, we have included the 8 models described below and summarised in Table 1. The arms in each model (except for contour maps) are given names for convenience.
  • Taylor_Cordes_1992: This model is bassed on radio and optical observations of H II regions. We use the model parameters presented in their Table 1 (Taylor & Cordes, 1993). There are four arms in this model (Arm1, Arm2, Arm3, Arm4).
  • Drimmel_NIR_2000: This model is based on Galactic plane emission profiles in the K band using COBE data. The model is publicly available. There are two arms (1_arm, 2_arm) and two inter-arm regions (3_interarm, 4_interarm) in this model.
  • Levine_2006: This model implements the logarithmic spiral framework as described in Levine et al. (2006). It is based on HI observations. We have adopted the model from their Table 1. There are four arms in this model (Arm1, Arm2, Arm3, Arm4).
  • Hou_Han_2014: This model is built upon the polynomial-logarithmic formulation introduced by Hou & Han (2014) and is based on a combination of H II, giant molecular clouds, and methanol MASER catalogs. Their model is publicly available. There are six arms in this model (Norma, Scutum-Centaurus, Sagittarius-Carina, Perseus, Local, Outer).
  • Reid_2019: This model is based on high precision radio astrometry of MASERS (Reid et al., 2019). We have adopted the model from their Table 2. There are seven arms in this model (3-kpc, Norma, Sct-Cen, Sgr-Car, Local, Perseus, Outer).
  • Poggio_2021: Using Gaia EDR3 astrometry of upper-main sequence stars, Poggio et al. (2021) generated 2D contour maps of spiral arms. Their data is available publicly, and also included in the package with their permission.
  • Gaia_2022: Using OB stars selected using Gaia Astrophysical parameters & DR3 astrometry, Gaia Collaboration et al. (2023) generated 2D contour maps of spiral arms. Their data is available publicly, and also included in the package with their permission.
  • Drimmel_Ceph_2024: Recently, new distances to 3425 classical cepheids were derived by Skowron et al. (2024) with a mean distance uncertainty of about 6%. From this Drimmel et al. (2024) used a subsample of dynamically young Milky Way Cepheids to construct models for spiral arms. Their model is publicly available but also included in the package as a userfriendly pickle file, with their permission. There are four arms in this model (Scutum, Sag-Car, Orion, Perseus).
This list is by no means exhaustive, but is instead a first attempt to compile spiral arms based on a variety of different tracers.
2 Example Usage & Scientific application
As an example of a science case, we reproduce figures from (Khanna et al., 2024, hereafter K24) where SpiralMap was used. In particular, K24 constructed a model for the stellar density distribution in Red Clump stars in the Milky Way, and then compared the residuals of their best-fit models with the locations of non-axisymmetric structures such as spiral arms. Figure 4 shows two of their residual plots overlaid with spiral models (Drimmel_NIR_2000, Drimmel_Ceph_2024, & Reid_2019) using SpiralMap.
Figure 4: An example science case where SpiralMap can be useful. Here we show two figures reproduced from K24 (with permission), comparing the residuals between a best-fit model and data, and the locations of spiral arm models in the Milky Way. Panel(a) shows the Drimmel_NIR_2000, Drimmel_Ceph_2024, & Reid_2019 overplotted together, while panel(b) shows only the Drimmel_NIR_2000 model. Both cases are shown in polar projection.
A very basic example of using SpiralMap is shown below, where we access all information about one particular arm (’Sag-Car’) in one particular model Drimmel_Ceph_2024,
##### Readout a single arm from a single model ###
import SpiralMap as sp
from SpiralMap import polar_style
Rsun=8.277
spirals = sp.main_(Rsun=Rsun)Ψ
use_model = ‘Drimmel_Ceph_2024’
spirals.getinfo(model=use_model)
plotattrs = {‘plot’:False}
spirals.readout(plotattrs,model=use_model,arm=‘Sag-Car’)
More common examples are shown in the accompanying Jupyter notebook.
SK acknowledges support from the European Union’s Horizon 2020 research and innovation program under the GaiaUnlimited project (grant agreement No 101004110). We thank Ronald Drimmel & Eloisa Poggio for useful suggestions. This work has used the following additional software products: Matplotlib (Hunter, 2007); IPython (Pérez & Granger, 2007); Pandas (pandas development team, 2020); Astropy, a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2018); NumPy (Harris et al., 2020).

### v2

@@
-SpiralMap: A Python library of the Milky Way’s spiral arms
+arXiv:2506.11383v2[astro-ph.IM] 01 Mar 2026
+arXiv:2506.11383v2[astro-ph.IM] 01 Mar 2026
 SpiralMap: A Python library of the Milky Way’s spiral arms
@@
 Over the years, various groups have surveyed the Galaxy across wavelengths (Radio to Optical) and have deduced the rich variety of spiral structure present in the disc.
-While some of the papers in literature provide machine readable data to trace out the spiral arms in their model, often this can be a cumbersome exercise for a user simply interested in extracting the coordinates and/or overplotting the spiral arms on another plot of interest, such as while comparing the locations of the arms to features in the velocity field (Khanna et al., 2023; Poggio et al., 2024, etc.).
+While some of the papers in literature provide machine readable data to trace out the spiral arms in their model, often this can be a cumbersome exercise for a user simply interested in extracting the coordinates and/or overplotting the spiral arms on another plot of interest, such as while comparing the locations of the arms to features in the velocity field (Khanna et al., 2023; Poggio et al., 2025, etc.).
 We show an example science application in Figure 4.
@@
 Taylor_Cordes_1992 Model based on HII (Taylor & Cordes, 1993).
+Vallee_1995 Model based on Galactic magnetic field/Dust/stars (Bower et al., 1995).
 Drimmel_NIR_2000 Model based on Galactic plane emission in the NIR (Drimmel, 2000).
@@
 Gaia_2022 Map based on OB stars (Gaia Collaboration et al., 2023)
-Drimmel_Ceph_2024 Model based on Cepheid variables (Drimmel et al., 2024).
+Drimmel_Ceph_2024 Model based on Cepheid variables (Drimmel et al., 2025).
 Figure 1: Cartesian projection of the Drimmel_Ceph_2024 model shown for a particular arm (Sag-Car).
@@
 We show the Taylor_Cordes_1992 & Poggio_2021 models in HC (a) and GC (b) frames, and similarly the Drimmel_NIR_2000 & Poggio_2021 models in HC (c) and GC (d) frames.
-1 Available models
-In the current version, we have included the 8 models described below and summarised in Table 1. The arms in each model (except for contour maps) are given names for convenience.
-  • Taylor_Cordes_1992: This model is bassed on radio and optical observations of H II regions.
+I Available models
+In the current version, we have included the 9 models described below and summarised in Table 1. The arms in each model (except for contour maps) are given names for convenience.
+  • Taylor_Cordes_1992: This model is based on radio and optical observations of H II regions.
 We use the model parameters presented in their Table 1 (Taylor & Cordes, 1993).
 There are four arms in this model (Arm1, Arm2, Arm3, Arm4).
+  • Vallee_1995: This model is based on a combination of the Galactic magnetic field, HI, HII, CO, and Dust distribution.
+We use their best model as depicted in their figure 3 (Bower et al., 1995).
+There are four arms in this model (Sagittarius, Scutum, Norma, Perseus).
   • Drimmel_NIR_2000: This model is based on Galactic plane emission profiles in the K band using COBE data.
@@
 Their model is publicly available.
-There are six arms in this model (Norma, Scutum-Centaurus, Sagittarius-Carina, Perseus, Local, Outer).
+There are six arms in this model (Arm1, Arm2, Arm3, Arm4, Arm5, Arm6).
+Additionally their local arm is included as a regular logarithmic spiral.
   • Reid_2019: This model is based on high precision radio astrometry of MASERS (Reid et al., 2019).
@@
 This list is by no means exhaustive, but is instead a first attempt to compile spiral arms based on a variety of different tracers.
-2 Example Usage & Scientific application
-As an example of a science case, we reproduce figures from (Khanna et al., 2024, hereafter K24) where SpiralMap was used.
+II Example Usage & Scientific application
+As an example of a science case, we reproduce figures from (Khanna et al., 2025, hereafter K24) where SpiralMap was used.
 In particular, K24 constructed a model for the stellar density distribution in Red Clump stars in the Milky Way, and then compared the residuals of their best-fit models with the locations of non-axisymmetric structures such as spiral arms.